# OmniVLM: A Token-Compressed, Sub-Billion-Parameter Vision-Language Model for Efficient On-Device Inference

### v1

@@ 2.1 Vision-Language Models: Evolution and Architectures
Building an instruction following agent has been an active field of study in computer vision. Vision language models have achieved impressive performance, since the introduction of LLaVA  [Liu et al., 2023b, a], a large multimodal model that combines a vision encoder and a language model to create an end-to-end trained system for general-purpose visual and language understanding. Their work generated multimodal instruction-following data using GPT-4, thereby addressing the scarcity of such data. Two evaluation benchmarks were introduced with diverse and challenging tasks to assess the model’s capabilities. Additionally, when fine-tuned on ScienceQA [Lu et al., 2022], a multimodal reasoning dataset, LLaVA in conjunction with GPT-4 achieves state-of-the-art accuracy. The research on vision language models has been further pushed by efforts in both academia and industries. Recent progresses include PaliGemma [Beyer et al., 2024] built on top of Gemma-2B [Team et al., 2024], Qwen2-VL [Bai et al., 2023, Wang et al., 2024a] built on top of QwenLM [Yang et al., 2024], Llama3.2-Vision built on top of Llama [Meta, 2024a, Dubey et al., 2024] and InternVL [Chen et al., 2024b] built from vision transformer [Dosovitskiy, 2020] with language models like QLLaMA [Cui et al., 2023] and Vicuna [Zheng et al., 2023] as middleware.
@@ 2.2 Edge Deployment and Inference Optimization
Due to constraints on memory and computing capability, deploying language models on edge devices such as smartphones and PC has been a challenge. Techniques such as quantization, pruning, and KV cache are widely used. The GGML library [Gerganov, 2023-2024a] provides necessary functions for fast inference of large language models and the gguf format for model weights, based on which abundant high-level inference code have been developed, such as llama.cpp [Gerganov, 2023-2024b] and ctransformers [Marella, 2023-2024]. Applications that are designed particularly for local runs have emerged in recent years. Notable softwares include Ollama [Ollama team, 2023-2024], MLC [MLC team, 2023-2024] and etc.

### v2

@@ 2.1 Vision-Language Models: Evolution and Architectures
Building an instruction following agent has been an active field of study in computer vision. Vision language models have achieved impressive performance, since the introduction of LLaVA  [Liu et al., 2023b, a], a large multimodal model that combines a vision encoder and a language model to create an end-to-end trained system for general-purpose visual and language understanding. Their work generated multimodal instruction-following data using GPT-4, thereby addressing the scarcity of such data. Two evaluation benchmarks were introduced with diverse and challenging tasks to assess the model’s capabilities. Additionally, when fine-tuned on ScienceQA [Lu et al., 2022], a multimodal reasoning dataset, LLaVA in conjunction with GPT-4 achieves state-of-the-art accuracy. The research on vision language models has been further pushed by efforts in both academia and industries. Recent progresses include PaliGemma [Beyer et al., 2024] built on top of Gemma-2B [Team et al., 2024], Qwen2-VL [Bai et al., 2023, Wang et al., 2024a] built on top of QwenLM [Yang et al., 2024], Llama3.2-Vision built on top of Llama [Meta, 2024a, Dubey et al., 2024] and InternVL [Chen et al., 2024c] built from vision transformer [Dosovitskiy, 2020] with language models like QLLaMA [Cui et al., 2023] and Vicuna [Zheng et al., 2023] as middleware.
@@ 2.2 Edge Deployment and Inference Optimization
Due to constraints on memory and computing capability, deploying language models on edge devices such as smartphones and PC has been a challenge. Techniques such as quantization, pruning, and KV cache are widely used. The GGML library [Gerganov, 2023-2024a] provides necessary functions for fast inference of large language models and the gguf format for model weights, based on which abundant high-level inference code have been developed, such as llama.cpp [Gerganov, 2023-2024b] and ctransformers [Marella, 2023-2024]. Applications [Chen and Li, 2024, Chen et al., 2024b] that are designed particularly for local runs have emerged in recent years. Notable softwares include Ollama [Ollama team, 2023-2024], MLC [MLC team, 2023-2024] and etc.